OpenStudy (anonymous):

given the equation below solve for x x^2-16x+64=2^2

OpenStudy (doggy):

please post this in the math part of OS

OpenStudy (anonymous):

x^2-16x+64=2^2 x^2-16x+64-4=0 x^2-16x+60=0 delta=256-240 delta=16 x=(16+-4)/2 x'=10 x''=6